# A Comprehensive Study on Large Language Models for Mutation Testing

### v4

@@ 2.2.1. Traditional Approaches
PIT: PIT (Coles et al., 2016) is a widely used mutation testing tool on the Java bytecode level, which supports 29 well-designed mutation operators and four groups of active mutation operator sets. In our study, we employ PIT 1.14.4 and leverage the ALL group which activates all operators.
Major: Major (Just et al., 2011) is another high-impact mutation testing framework, that supports nine operators. In our study, we use Major 1.3.5 and involve all mutation operators.
LEAM: LEAM (Tian et al., 2022, [n. d.]) is the state-of-the-art custom-trained-model-based approach, which employs sequence-to-sequence models to generate mutations. Its model is trained on a large corpus of real-world bugs and their corresponding patches, allowing it to learn patterns for generating more effective mutations.
BERT: BERT (Degiovanni and Papadakis, 2022) is the state-of-the-art mutation generation approaches based on the small-scale pre-trained model, BERT (Devlin et al., 2018). Equipped with the pre-trained model, BERT transforms the task of mutation generation to code completion under the surrounding context.
@@ 2.2.2. LLM-based Approaches
Our study incorporates two existing LLM-based mutation generation approaches (i.e., BugFarm (Ibrahimzada et al., 2023) and LLMorpheus (Tip et al., 2025)), and introduces a new prompt design for mutation generation, which will be presented later.
Existing LLM-based Approaches: We adopt two the state-of-the-art LLM-based mutation generation approaches, BugFarm (Ibrahimzada et al., 2023) and LLMorpheus (Tip et al., 2025). BugFarm trains a small-scale model for selecting lines for mutation, and then utilizes LLMs to mutate the target line. LLMorpheus performs mutation given the whole method as context. Note that LLMorpheus was originally designed for JavaScript. Upon examining its prompt, we found it to be language-agnostic. Therefore, we reimplemented the code-extraction component for Java (such as parsing ASTs and extracting the code context for mutation) while retaining the original prompt. For both approaches, we employ GPT-4o (Achiam et al., 2023), the current most widely used model, as their backend LLM.
@@
LLMs with Our Prompt: Different from the existing LLM-based approaches, we propose a new prompt engineering for mutation generation, shown as Figure 1. For ease of presentation, in the following part of this paper, we call our prompt LLMut.
@@
To craft effective prompts, we follow the best practices, which suggest that prompts should comprise four aspects: Instruction, Context, Input Data, and Output Indicator (Guo et al., 2024; Ma et al., 2024). In the Instructions, we direct the LLMs to generate mutants for the target code element. The placeholder MUT_NUMBER specifies the number of mutations required, while the placeholder CODE_ELEMENT specifies the target code snippet to be mutated. In the Context, we clearly state that a mutant is the concept of mutation testing, and additionally provide various sources of information to observe the impact on the performance of LLMs, including the whole Java method surrounding the target code element and few-shot examples sampled from real-world bugs from another benchmark. In the Input Data, we specify the code element to be mutated and the number of mutants to generate. In the Output Indicator, we specify the JSON file format for mutation outputs, facilitating further experiments. In addition, our prompts need few-shot examples, which should originate from real bugs, enabling LLMs to learn how to mutate code from real bugs. We further explore the different prompt settings in Section 4.1 and the impact of prompt turbulence in Section 4.2.
@@
To avoid the few-shot examples leaking information, beyond the evaluation datasets (i.e., as shown later, they are Defects4J and ConDefects), we employ another benchmark, QuixBugs (Lin et al., 2017), which comprises 40 real-world Java bugs and is commonly used by the testing and debugging community (Zhu et al., 2023; Jiang et al., 2023; Xia and Zhang, 2022; Li et al., 2023b). Referring to the study by Deng et al. (Deng et al., 2024), LLMs perform best in generating valid code when using around six few-shot examples. Therefore, we sample six bugs from QuixBugs for few-shot learning in our prompt. To guarantee the diversity of the examples, we randomly select one bug from the dataset and assess whether its modification pattern is similar to the examples already collected. If it is different, we add it to our collection and continue, until we have collected all six examples, shown as Table 2. As LLM capabilities continue to advance, the optimal number of few-shot examples may vary. We therefore investigate the impact of different numbers of few-shot examples in Section 4.1.3.
@@
To determine the number of mutations generated in a prompt, we adopt the setting of PIT (Coles et al., 2016). We use PIT-generated mutations with the universal set of operators and find that PIT generates 1.15 mutations per line of code. Therefore, based on the given code context, we generate one mutation per line. We explore the influence of context length on mutation generation in Section 4. Additionally, for the temperature, we use the default value for each LLM.
Compared with the prompts used in BugFarm222https://github.com/Intelligent-CAT-Lab/BugFarm and LLMorpheus333https://github.com/githubnext/llmorpheus, our prompt design offers several advantages:
it incorporates real-world bugs as few-shot examples, which have been shown to significantly enhance LLM performance in many code-related tasks;
it determines the number of generated mutants based on the size of the target code, rather than on a fixed number;
it specifies the mutation location using the target statement itself rather than line numbers, enabling more precise localization (Fan et al., 2023a); and
it employs JSON output formatting, which is widely used in LLM fine-tuning and downstream tasks.
Based on LLMut, our study comprehensively covers the most widely studied LLMs. Specifically, we select a representative set of LLMs with code capabilities, both from commercial closed-source and open-source models. Following recent studies on evaluating the code capabilities of LLMs (Wang et al., 2024; Guo et al., 2024; Yuan et al., 2024), we adopt the closed-source LLMs from the GPT family, including GPT-3.5 (GPT-3.5-Turbo) (Achiam et al., 2023), GPT-4o (Achiam et al., 2023), and GPT-4o-M (GPT-4o-Mini) (Achiam et al., 2023)), and open-source LLMs, including SC-16b (StarChat--16b), CL-13b (CodeLlama-Instruct-13b), DS-236b (DeepSeek-Coder-V2-236b), and DS-671b (DeepSeek-V3-671b). The details of these models are shown as Table 1. Note that OpenAI has not officially released the parameter size of their closed-source models. The parameter size of the open-sourced models we studied ranges from 13 billion to 671 billion. Moreover, the selected LLMs represent a balance between general-purpose models (e.g., GPT-4o and DeepSeek-V3-671b), which are trained for a wide range of natural language and reasoning tasks, and code-specialized models (e.g., CodeLlama-Instruct-13b and DeepSeek-V2-Coder-236b), which are further fine-tuned on large-scale code corpora. This selection allows us to evaluate mutation generation across both broadly capable models and those explicitly optimized for programming tasks, ensuring the conclusions are representative of the current landscape of LLMs.
@@ 2.3. Datasets
As one of the major tasks of our study is to evaluate the behavioral similarity between mutations and real-world bugs, we selected datasets that include real-world bugs along with their corresponding fault-revealing test cases. Specifically, we employ the datasets with the following properties:
@@
The bugs of the datasets should be real-world bugs so that we can compare the difference between mutations and real bugs.
@@
Every bug is accompanied by at least one bug-triggering test because we need to measure whether the mutations affect the execution of bug-triggering tests.
To this end, we employ the Defects4J 2.0 (Just et al., 2014b) and ConDefects (Wu et al., 2024) to evaluate the mutation generation approaches, shown as Table 3. In total, we conducted experiments on 851 bugs.
Defects4J 2.0 is a widely used benchmark in the field of mutation testing (Ojdanic et al., 2023; Just et al., 2014c; Kaufman et al., 2022; Degiovanni and Papadakis, 2022; Tian et al., 2022; Kim et al., 2022), which contains historical bugs from open-source projects of diverse domains, ensuring a broad representation of real-world bugs. We select the most widely used 12 projects from Defects4J 2.0, which contain 605 bugs in total. However, from Table 1 and Table 3, we observe that the periods of the Defects4J bugs are earlier than the LLMs’ training time, which may introduce data leakage. Therefore, we supplement another dataset, ConDefects (Wu et al., 2024), designed to address the data leakage concerns. ConDefects consists of tasks from AtCoder444https://atcoder.jp programming contest. To prevent data leakage, we use data from the four most recent months following the release of the LLMs (i.e., from March 1, 2024, to June 30, 2024), during which we collected 246 Java bugs. Among them, only the most recent model, DeepSeek-V3-731b, cannot guarantee data leakage-free, since it is released after the ConDefects dataset.
@@ 2.4. Evaluation Metrics
In this subsection, we present the evaluation metrics used in our study to assess the quality of the generated mutations comprehensively. We evaluate the mutations from multiple perspectives, as mentioned before. Note that only relying on the mutation score is insufficient for a meaningful assessment. A higher mutation score does not necessarily indicate better quality mutations, as it only reflects the proportion of killed mutants rather than their effectiveness. However, we still report the mutation score of each mutation generation approach later in Section 3.
To better understand the metrics, we first illustrate the categories of mutations. Let all generated mutations be the set , where LLMs may generate syntactically incorrect mutations that can not pass compilation. Thus, we denote the compilable mutations as the set . Note that the set contains two types of mutations that should be removed, the duplicate mutations (denoted as ) and equivalent mutations (denoted as ). Therefore, we have . The duplicate mutations refer to those identical to the original code or other mutations. In mutation testing, the ideal set of mutations to employ is from the set , excluding equivalent mutations that have no contribution to discovering the weakness of tests. However, determining the equivalent mutations set is undecidable, so we follow the common practice and use the set as a practical approximation to perform mutation testing (Diniz et al., 2021; Tian et al., 2022; Zhang et al., 2022; Wang et al., 2025). Therefore, let denote the set of killed mutants among the compilable and non-duplicate ones (i.e., ). The mutation score is then computed as follows:
Note that traditional mutation generation approaches, such as Major and PIT, produce very few duplicate mutations (i.e., is 0). Hence, early studies (Kintis et al., 2017) defined equivalent mutations as those semantically equivalent to the original code, which may implicitly include , which is an empty set. However, LLM-based approaches tend to generate a large number of non-mutated duplicates. Moreover, while duplicate mutations can be easily filtered, identifying equivalent mutations remains challenging (Tian et al., 2024). Therefore, in this paper, we refine the definition of equivalent mutations to be those that are syntactically different but semantically equivalent to the original code.
@@ 2.4.1. RQ1: Effectiveness
Evaluating the effectiveness of mutation generation approaches requires assessing how well the generated mutations mimic real bugs. A key aspect of this evaluation is understanding whether LLM-generated mutations exhibit fault characteristics that align with real-world defects.
@@
To quantify this effectiveness, we employ multiple metrics that capture different aspects of behavioral similarity between mutations and real bugs. These metrics assess whether bug-triggering tests can detect LLM-generated mutations (i.e., RQ1.1), how closely mutations resemble real faults based on test case failures (i.e., RQ1.2), and the diversity of syntactic changes introduced by different mutation generation techniques (i.e., RQ1.3). The metrics of these sub-RQs are listed below.
@@
RQ1.1: Real Bug Detection Rate: Following the approach of Just et al. (Just et al., 2014c), we evaluate the effectiveness of mutation generation by measuring its ability to detect real bugs using developer-written tests. For a given real bug (which may span multiple lines), we apply a mutation generation approach to create a set of mutations on the fixed version of the program. If the test cases that kill a mutation also fail on the real bug, we consider that the mutation generation approach has successfully “detected” this real bug. To quantify this effectiveness, we compute the Real Bug Detection Rate, which represents the proportion of real bugs for which at least one generated mutation is detected by the same test cases that reveal the real bug. This metric provides insight into how well LLM-generated mutations align with actual software defects in a testing context.
@@
RQ1.2: Average Ochiai Coefficient: This metric estimates the semantic similarity between a mutation and its corresponding real bug based on their passing and failing test cases (Offutt and Untch, 2001). Ochiai coefficient measures how a mutation resembles its corresponding real bug. The Ochiai coefficient was originally introduced as a general similarity measure but later became widely adopted in spectrum-based fault localization (SBFL) (Zou et al., 2019). In this study, we use the Ochiai coefficient in its original form to assess how closely a mutation resembles its corresponding real bug, which is consistent with the study of Ojdanic et al. (Ojdanic et al., 2023).
@@
More concretely, let and be a mutation and its corresponding real bug, and let and be the set of their failing test cases, respectively. The coefficient of is computed by the following formula:
@@
Therefore, if the Ochiai coefficient of is , it is semantically equal to its corresponding bug . Given a bug , its Ochiai coefficient is calculated as the mean Ochiai coefficient across all its corresponding de-duplicated mutations (i.e., ), shown as the following formula:
@@
Given a mutation generation approach and a set of bugs , the Average Ochiai Coefficient of the approach is calculated as the mean Ochiai coefficient across all bugs, shown as the following formula:
@@
RQ1.2: High Ochiai Bug Count: Following the study (Ojdanic et al., 2023), a bug and a set of mutations are considered to have high Ochiai if they satisfy . This threshold indicates a strong semantic similarity between the generated mutations and the real bug. For each mutation generation approach, we compute the number of bugs that meet this criterion.
RQ1.2: Coupling Rate: Given a real bug, one mutant is considered coupling if the set of test cases that kill the mutant overlaps with the bug-triggering tests. Note that each bug we analyzed has at least one failing test to reveal the bug, called bug-triggering test. The Coupling Rate refers to the proportion of coupled mutants within the set of de-duplicated mutations (i.e., ). Note that Real Bug Detection Rate and Coupling Rate measure effectiveness from different perspectives. Real Bug Detection Rate evaluates effectiveness at the bug level, i.e., the proportion of bugs that can be revealed by at least one mutant. In contrast, Coupling Rate evaluates effectiveness at the mutation level, i.e., the proportion of mutants whose failing tests overlap with bug-triggering tests. For instance, a mutation approach may generate a large number of high-quality mutants for only a few specific bugs, but few mutants for others. This can result in a high Coupling Rate (since the generated mutants are strongly aligned with the bug-triggering tests of those few bugs) but a low Real Bug Detection Rate (since many other bugs remain undetected). Conversely, another approach may generate mutants that cover a wide range of bugs, thereby achieving a high Real Bug Detection Rate. However, if the mutants for each bug exhibit only weak coupling with the bug-triggering tests, the overall Coupling Rate will remain low. Therefore, these two metrics are complementary, with each reflecting only one aspect of mutation effectiveness.
@@
RQ1.2: Coupling Categorization: We categorize each generated mutation into distinct coupling scales based on their similarity to real bugs, following the classification proposed by (Gay and Salahirad, 2023), and analyze their distribution across different approaches. Specifically, given a compilable mutation (i.e., an element from the set ) on the fixed version of a real bug (which can be uncovered by at least one triggering test), we can classify it into the following six scales based on its behavior:
@@
Strong Substitution: All triggering tests fail and no additional tests fail, which represents the mutation is a semantic replacement of the real bug under the given test suite.
Strong Substitution + Additional Tests Fail: All triggering tests fail but additional non-triggering tests fail.
@@
Partial Substitution + Additional Tests Fail: Some triggering tests fail and some additional non-triggering tests fail.
@@
Not Detected: No tests fail (i.e., the mutation is alive under the given test suite).
Later, we will classify the mutations of each approach and analyze their distributions.
RQ1.3: AST Node Diversity: We prefer the set of mutations with diverse syntactic forms, i.e., method invocations, field accesses, etc. Therefore, we parse the code before and after mutation, and check the diversity of newly introduced AST nodes. In other words, we identify the unique AST node types introduced by each approach’s generated mutants, relative to the set of node types present in the original program. The approach with a greater variety of AST node types is considered superior, as it indicates a broader exploration of the syntactic space and enhances the likelihood of detecting weaknesses in the test suite.
RQ1.3: AST Edit Distance: The AST edit distance measures the structural difference between the original and mutated code by quantifying the number of tree transformation operations (i.e., insertions, deletions, and modifications) needed to convert one AST into another. This metric evaluates the extent of syntactic changes introduced by mutations. A higher edit distance suggests more significant syntactic changes. The AST Edit Distance metric is well-suited for structured code, as it captures syntactic differences at the tree level rather than merely comparing token sequences, as in BLEU and similar metrics. However, it still has inherent limitations, such as its inability to fully reflect the impact of mutations on program behavior. We further discuss these limitations later in Section 5.

### v5

@@ 2.2.1. Traditional Approaches
PIT: PIT (Coles et al., 2016) is a widely used mutation testing tool on the Java bytecode level, which supports 29 well-designed mutation operators and four groups of active mutation operator sets. In our study, we employ PIT 1.14.4 and leverage the ALL group, which activates all operators.
Major: Major (Just et al., 2011) is another high-impact mutation testing framework that supports nine operators. In our study, we use Major 1.3.5 and involve all mutation operators.
LEAM: LEAM (Tian et al., 2022, [n. d.]) is the state-of-the-art custom-trained-model-based approach, which employs sequence-to-sequence models to generate mutants. Its model is trained on a large corpus of real-world bugs and their corresponding patches, allowing it to learn patterns for generating more effective mutants.
BERT: BERT (Degiovanni and Papadakis, 2022) is the state-of-the-art mutant generation approaches based on the small-scale pre-trained model, BERT (Devlin et al., 2018). Equipped with the pre-trained model, BERT transforms the task of mutant generation to code completion under the surrounding context.
@@ 2.2.2. LLM-based Approaches
Our study incorporates two existing LLM-based mutant generation approaches (i.e., BugFarm (Ibrahimzada et al., 2025) and LLMorpheus (Tip et al., 2025)), and introduces a new prompt design for mutant generation, which will be presented later.
Existing LLM-based Approaches: We adopt two the state-of-the-art LLM-based mutant generation approaches, BugFarm (Ibrahimzada et al., 2025) and LLMorpheus (Tip et al., 2025).
BugFarm (Ibrahimzada et al., 2025) trains a small-scale model for selecting lines for mutating, and then utilizes LLMs to mutate the target line. Based on the prompt template given in the repository of BugFarm222https://github.com/Intelligent-CAT-Lab/BugFarm, we can find that BugFarm numbers all lines of a Java method and uses a lightweight model to predict the target line numbers to mutate. These line numbers are then provided to the LLM. BugFarm then instructs the LLM to generate three different mutants for the Java method by changing the specified lines. Finally, BugFarm outputs each mutant as a complete mutated method, with the mutation embedded in the full method.
LLMorpheus (Tip et al., 2025) is an LLM-based mutation testing approach for JavaScript. Upon examining its prompt from its repository333https://github.com/githubnext/llmorpheus, we found it to be language-agnostic. Therefore, we reimplement the code-extraction component for Java (such as parsing ASTs and extracting the code context for mutants) while retaining its original prompt. When porting LLMorpheus to Java, we retain its default configuration. LLMorpheus masks specific syntactic elements (e.g., the conditional expression of an if statement) with a PLACEHOLDER to mark the target mutation location, and instructs the LLM to fill in this placeholder. For each masked location, LLMorpheus generates three mutants. By default, LLMorpheus provides up to 200 lines of surrounding context and instructs the LLM to return only the code fragment that should replace the placeholder.
In our study, both BugFarm and LLMorpheus are equipped with GPT-4o (Achiam et al., 2023) and DeepSeek-V3-671B (Liu et al., 2024), representing the current leading closed- and open-source LLMs, respectively.
@@
LLMs with Our New Prompt: Different from the existing LLM-based approaches, we propose a new prompt engineering for mutant generation, shown as Figure 1. For ease of presentation, in the following part of this paper, we call our prompt LLMut.
@@
To craft effective prompts, we follow the best practices which suggest that prompts should comprise four aspects: Instruction, Context, Input Data, and Output Indicator (Guo et al., 2024; Ma et al., 2024). In the Instructions, we direct the LLMs to generate mutants for the target code element. The placeholder MUT_NUMBER specifies the number of mutants required, while the placeholder CODE_ELEMENT specifies the target code snippet to be mutated. In the Context, we clearly state that a mutant is the concept of mutation testing, and additionally provide various sources of information to observe the impact on the performance of LLMs, including the whole Java method surrounding the target code element and few-shot examples sampled from real-world bugs from another benchmark. In the Input Data, we specify the code element to be mutated and the number of mutants to generate. In the Output Indicator, we specify the JSON file format for outputs, facilitating further experiments. In addition, our prompts need few-shot examples, which should originate from real bugs, enabling LLMs to learn how to mutate code from real bugs. We further explore the different prompt settings in Section 4.1.
@@
To avoid the few-shot examples leaking information, beyond the evaluation datasets (i.e., as shown later, they are Defects4J and ConDefects), we employ another benchmark, QuixBugs (Lin et al., 2017), which comprises 40 real-world Java bugs and is commonly used by the testing and debugging community (Zhu et al., 2023; Jiang et al., 2023; Xia and Zhang, 2022; Li et al., 2023b). As shown later, our experimental results indicate that providing six examples results in the best overall performance, consistent with observations by Deng et al. (Deng et al., 2024). Therefore, we sample six bugs from QuixBugs for few-shot learning in our prompt. To guarantee the diversity of the examples, we randomly select one bug from the dataset and assess whether its modification pattern is similar to the examples already collected. If it is different, we add it to our collection and continue until we have collected all examples, shown as Table 2. As LLM capabilities continue to advance, the optimal number of few-shot examples may vary. We therefore investigate the impact of different numbers of few-shot examples in Section 4.1.3. Note that adopting few-shot examples may also have drawbacks, which we will discuss later.
@@
To determine the number of mutants generated in a prompt, we adopt the setting used in PIT (Coles et al., 2016). We use PIT-generated mutants with the universal set of operators and find that PIT generates 1.15 mutants per line of code. Therefore, based on the given code context, we generate one mutant per line. We explore the influence of context length on mutant generation in Section 4. Additionally, for the temperature, we use the default value for each LLM.
Each LLM-based approach in our study adopts a different prompt-engineering strategy. For example, LLMut incorporates few-shot examples and uses a structured JSON output format, whereas LLMorpheus masks high-value code fragments and instructs the LLM to fill in the masked content to ensure the quality of the output mutants. Note that the different prompt-engineering strategies are complementary, with different approaches emphasizing different aspects of the mutants.
Based on LLMut, our study comprehensively covers the most widely studied LLMs. Specifically, we select a representative set of LLMs with code capabilities, both from commercial closed-source and open-source models. Following recent studies on evaluating the code capabilities of LLMs (Wang et al., 2024; Guo et al., 2024; Yuan et al., 2024), we adopt the closed-source LLMs from the GPT family, including GPT-3.5-Turbo (GPT-3.5) (Achiam et al., 2023), GPT-4o (Achiam et al., 2023), and GPT-4o-Mini (GPT-4o-M) (Achiam et al., 2023)), and open-source LLMs, including StarChat--16b (SC-16b), CodeLlama-Instruct-13b (CL-13b), DeepSeek-Coder-V2-236b (DS-236b), and DeepSeek-V3-671b (DS-671b). The details of these models are shown as Table 1. Note that OpenAI has not officially released the parameter size of their closed-source models. The parameter size of the open-sourced models we studied ranges from 13 billion to 671 billion. Moreover, the selected LLMs represent a balance between general-purpose models (e.g., GPT-4o and DeepSeek-V3-671b), which are trained for a wide range of natural language and reasoning tasks, and code-specialized models (e.g., CodeLlama-Instruct-13b and DeepSeek-V2-Coder-236b), which are further fine-tuned on large-scale code corpora. This selection allows us to evaluate mutant generation across both broadly capable models and those explicitly optimized for programming tasks, ensuring the conclusions are representative of the current landscape of LLMs.
@@ 2.3. Datasets
As one of the major tasks of our study is to evaluate the behavioral similarity between mutants and real-world bugs, we selected datasets that include real-world bugs along with their corresponding fault-revealing test cases. Specifically, we employ the datasets with the following properties:
@@
The bugs of the datasets should be real-world bugs so that we can compare the difference between mutants and real bugs.
@@
Every bug is accompanied by at least one bug-triggering test because we need to measure whether the mutants affect the execution of bug-triggering tests.
To this end, we employ the Defects4J 2.0 (Just et al., 2014b) and ConDefects (Wu et al., 2024) to evaluate the mutant generation approaches, shown as Table 3. In total, we conducted experiments on 851 bugs.
Defects4J 2.0 is a widely used benchmark in the field of mutation testing (Ojdanic et al., 2023; Just et al., 2014c; Kaufman et al., 2022; Degiovanni and Papadakis, 2022; Tian et al., 2022; Kim et al., 2022), which contains historical bugs from open-source projects of diverse domains, ensuring a broad representation of real-world bugs. We select the most widely used 12 projects from Defects4J 2.0, which contain 605 bugs in total. However, from Table 1 and Table 3, we observe that the periods of the Defects4J bugs are earlier than the LLMs’ training time, which may introduce data leakage. Therefore, we supplement another dataset, ConDefects (Wu et al., 2024), designed to address the data leakage concerns. ConDefects consists of tasks from AtCoder444https://atcoder.jp programming contest. To prevent data leakage, we use data from the four most recent months following the release of the LLMs (i.e., from March 1, 2024, to June 30, 2024), during which we collected 246 Java bugs. Among them, only the most recent model, DeepSeek-V3-731b, cannot guarantee data leakage-free, since it was released after the ConDefects dataset.
@@ 2.4. Evaluation Metrics
In this subsection, we present the evaluation metrics used in our study to comprehensively assess the quality of the generated mutants. We evaluate the mutants from multiple perspectives, as mentioned before. Note that only relying on the mutation score is insufficient for a meaningful assessment. A higher mutation score does not necessarily indicate better quality mutants, as it only reflects the proportion of killed mutants rather than their effectiveness. However, we still report the mutation score of each mutant generation approach later in Section 3.
To better understand the metrics, we first illustrate the categories of mutants. Let all generated mutants be the set , where LLMs may generate syntactically incorrect mutants that can not pass compilation. Thus, we denote the compilable mutants as the set . Note that the set contains two types of mutants that should be removed, the duplicate mutants (denoted as ) and equivalent mutants (denoted as ). Therefore, we have . The duplicate mutants refer to those identical to the original code or other mutants. Ideally, we should perform mutation testing against the mutant set of , excluding all duplicated mutants () and equivalent mutants (), which have no contribution to discovering the weakness of tests. Based on our definition, the duplicated mutants are syntactically identical mutants that are trivially decidable. However, determining the equivalent mutants set is undecidable, and is negligible compared to the universal mutant set (i.e., ). Therefore, we use the set of mutants as an approximation to compute the mutation score (i.e., ), which is the common practice of existing studies (Diniz et al., 2021; Tian et al., 2022; Zhang et al., 2022; Wang et al., 2025a). Therefore, let denote the set of killed mutants among the compilable and non-duplicate ones (i.e., ). The mutation score is then computed as follows:
Note that traditional mutant generation approaches, such as Major and PIT, produce very few duplicate mutants (i.e., is 0). Hence, early studies (Kintis et al., 2017) defined equivalent mutants as those semantically equivalent to the original code, which may implicitly include , which is an empty set. However, LLM-based approaches tend to generate a large number of non-mutated duplicates. Moreover, while duplicate mutants can be easily filtered, identifying equivalent mutants remains challenging (Tian et al., 2024). Therefore, in this paper, we refine the definition of equivalent mutants to be those that are syntactically different but semantically equivalent to the original code.
@@ 2.4.1. RQ1: Effectiveness
Evaluating the effectiveness of mutant generation approaches requires assessing how well the generated mutants mimic real bugs. A key aspect of this evaluation is understanding whether LLM-generated mutants exhibit fault characteristics that align with real-world defects.
@@
To quantify this effectiveness, we employ multiple metrics that capture different aspects of behavioral similarity between mutants and real bugs. These metrics assess whether bug-triggering tests can detect LLM-generated mutants (i.e., RQ1.1), how closely mutants resemble real faults based on test case failures (i.e., RQ1.2), and the diversity of syntactic changes introduced by different mutant generation techniques (i.e., RQ1.3). The metrics of these sub-RQs are listed below.
@@
RQ1.1: Real Bug Detection Rate: Following the approach of Just et al. (Just et al., 2014c), we evaluate the effectiveness of mutant generation by measuring its ability to detect real bugs using developer-written tests. For a given real bug (which may span multiple lines), we use a mutant-generation approach to create a set of mutants on the fixed version of the program. If the test cases that kill a mutant also fail on the real bug, we consider that the mutant generation approach has successfully “detected” this real bug. To quantify this effectiveness, we compute the Real Bug Detection Rate, which represents the proportion of real bugs for which at least one generated mutant is detected by the same test cases that reveal the real bug. In other words, it measures whether the generated mutants can reproduce the failing behaviors of a set of real bugs under tests. This metric provides insight into how well LLM-generated mutants align with actual software defects in a testing context.
@@
RQ1.2: Average Ochiai Coefficient: This metric estimates the semantic similarity between a mutant and its corresponding real bug based on their passing and failing test cases (Offutt and Untch, 2001). Ochiai coefficient measures how a mutation resembles its corresponding real bug. The Ochiai coefficient was originally introduced as a general similarity measure but later became widely adopted in spectrum-based fault localization (SBFL) (Zou et al., 2019). In this study, we use the Ochiai coefficient in its original form to assess how closely a mutant resembles its corresponding real bug, which is consistent with the findings of Ojdanic et al. (Ojdanic et al., 2023).
@@
More concretely, let and be a mutant and its corresponding real bug, and let and be the set of their failing test cases, respectively. The coefficient of is computed by the following formula:
@@
Therefore, if the Ochiai coefficient of is , it is semantically equal to its corresponding bug . Given a bug , the Ochiai coefficient for is calculated as the mean Ochiai coefficient across all its corresponding de-duplicated mutants (i.e., ), shown as the following formula:
@@
Given a mutant generation approach and a set of bugs , the Average Ochiai Coefficient of the approach is calculated as the mean Ochiai coefficient across all bugs, shown as the following formula:
@@
RQ1.2: High Ochiai Bug Count: Following the study (Ojdanic et al., 2023), a bug and a set of mutants are considered to have high Ochiai if they satisfy . This threshold indicates a strong semantic similarity between the generated mutants and the real bug. For each mutant generation approach, we compute the number of bugs that meet this criterion.
RQ1.2: Coupling Rate: Given a real bug, one mutant is considered coupling if the set of test cases that kill the mutant overlaps with the bug-triggering tests. Note that each bug we analyzed has at least one failing test to reveal the bug, called bug-triggering test. The Coupling Rate refers to the proportion of coupled mutants within the set of de-duplicated mutants (i.e., ).
Note that Real Bug Detection Rate and Coupling Rate assess effectiveness from different granularities. Real Bug Detection Rate evaluates effectiveness at the bug level—the proportion of bugs for which at least one generated mutant reproduces the failing behavior. Coupling Rate evaluates effectiveness at the mutant level—the proportion of mutants whose failing tests overlap with the bug-triggering tests. These two metrics capture different phenomena. For example, an approach may generate many high-quality mutants for only a few bugs while producing very few mutants for the remaining bugs. In this case, the Coupling Rate can be high (because many mutants align well with the few covered bugs), but the Real Bug Detection Rate remains low (because many bugs never receive any useful mutants). Conversely, an approach might generate mutants that touch a wide range of bugs, achieving a high Real Bug Detection Rate. However, if per-bug mutants only weakly reflect the bug-triggering behavior, the overall Coupling Rate may remain low. Therefore, the two metrics are complementary: Real Bug Detection Rate measures coverage across bugs, while Coupling Rate measures behavioral precision within mutants. Neither metric alone can fully represent the effectiveness of mutation generation.
@@
RQ1.2: Coupling Categorization: We categorize each generated mutant into distinct coupling scales based on their similarity to real bugs, following the classification proposed by (Gay and Salahirad, 2023), and analyze their distribution across different approaches. Specifically, given a compilable mutant (i.e., an element from the set ) on the fixed version of a real bug (which can be uncovered by at least one triggering test), we can classify it into the following six scales based on its behavior:
@@
Strong Substitution: All triggering tests fail and no additional tests fail, which represents the mutant is a semantic replacement of the real bug under the given test suite.
Strong Substitution + Additional Tests Fail: All triggering tests fail, but additional non-triggering tests fail.
@@
Partial Substitution + Additional Tests Fail: Some triggering tests fail, and some additional non-triggering tests fail.
@@
Not Detected: No tests fail (i.e., the mutant is surviving under the given test suite).
Later, we will classify the mutants of each approach and analyze their distributions.
RQ1.3: AST Node Diversity: We prefer the set of mutants with diverse syntactic forms, i.e., method invocations, field accesses, etc. Therefore, we parse the code before and after mutation, and check the diversity of newly introduced AST nodes. In other words, we identify the unique AST node types introduced by each approach’s generated mutants, relative to the set of node types present in the original program. The approach with a greater variety of AST node types is considered superior, as it indicates a broader exploration of the syntactic space and enhances the likelihood of detecting weaknesses in the test suite.
RQ1.3: AST Edit Distance: The AST edit distance measures the structural difference between the original and mutated code by quantifying the number of tree transformation operations (i.e., insertions, deletions, and modifications) needed to convert one AST into another. This metric evaluates the extent of syntactic changes introduced by mutations. A higher edit distance suggests more significant syntactic changes. The AST Edit Distance metric is well-suited for structured code, as it captures syntactic differences at the tree level rather than merely comparing token sequences, as in BLEU and similar metrics. However, it still has inherent limitations, such as its inability to fully reflect the impact of mutants on program behavior. We further discuss these limitations later in Section 5.